Becky would like to have at least $250,000 saved for her daughter’s college education. If she invests $80,000 in an education account paying 6.15% interest compounded quarterly, will she reach her goal in 18 years? Show all work to justify your answer and include appropriate units.


Answer by ikleyn(52800)

.

f = p * (1 + r) ^ n


f is the future value
p is the present value
r is the interest rate per time period
n is the number of time periods.


in your problem:


p = 80000
f = what you want to find
r = 6.15% interest per year compounded quarterly = 0.0615/4 per quarter
n = 4*18 = 72 compounding periods


formula becomes f = 80000 * (1 + 0.0615/4)^72


Calculate to get f = 239991.33.